OpenStudy (anonymous):

A ball is thrown upward and outward from the top edged of a 50 ft building it reaches its highest point 20 ft above and 10 ft out from the building. How far from the building is the ball when it hits the ground ?

OpenStudy (anonymous):

Are you familiar with the motion formulas?

OpenStudy (anonymous):

its a physics question...

OpenStudy (anonymous):

We are instructed to do it in analytic geometry

OpenStudy (anonymous):

mathematical modelling

OpenStudy (anonymous):

so it is an application of a parabola

OpenStudy (anonymous):

Can I just see your imagination in this problem. I can't imagine it Im noob at applications

OpenStudy (anonymous):

I'm solving in the view of physics. Though now you said it, it's probably more wise to use parabola to solve it.

OpenStudy (anonymous):

So, you have a parabola, with vertex(10,20) and a point on the parabola, (0,50)

OpenStudy (anonymous):

It is stated in the book that the answer must be 28.7 feet

OpenStudy (anonymous):

Should it be (60,70) ?

OpenStudy (anonymous):

Yeah, it should.

OpenStudy (anonymous):

I meant (10 , 70)

OpenStudy (anonymous):

Yeah...so the vertex is (10,70) there's a point(0,50), and we're trying to find the x-intercept. Sub all these into the parabola, \((x+k)^2 = 4a(y+h)\) What did you get?

OpenStudy (anonymous):

should it be ? (X - H)^2 = 4a(y-k)

OpenStudy (anonymous):

oh. Your convention is negative.

OpenStudy (anonymous):

I meant - 4a because its downward right ?

OpenStudy (anonymous):

But this is what confuses me how do you that that's a vertex :(

OpenStudy (anonymous):

Yup.Though that will be calc'ed later. because it's the highest point of the whole journey. the parabola is \((x-10)^2=-5(y-70)\) Solving with y=0 gives 28.7.

OpenStudy (anonymous):

so 4a = - 5 my god thank you Shadowys

OpenStudy (anonymous):

That was done by subbing (0,50) into it. You're welcome :)

OpenStudy (anonymous):

wait so it gives a quadratic equation ?

OpenStudy (anonymous):

Yup, it does.

OpenStudy (anonymous):

x^2 - 20x - 450?

OpenStudy (anonymous):

The answer is not exact -_-

OpenStudy (anonymous):

Um, how did you get that equation? from \((x−10)^2=−5(y−70)\)?

OpenStudy (anonymous):

That is the parabola of the thrown ball.

OpenStudy (anonymous):

X^2 - 20x +100 = -5( 0 - 70)

OpenStudy (anonymous):

X^2 - 20x - 250

OpenStudy (anonymous):

it's 28.708289 to be exact. Yes, that gives the correct answer. Though you could save some steps by not expanding.

OpenStudy (anonymous):

Aw how did you do it :O

OpenStudy (anonymous):

let me guess square root ?

OpenStudy (anonymous):

Now I get I FORGOT THE BASICS HAHAHAHAH I forgot to square both sides . Thanks again Shadowy! :)

OpenStudy (amistre64):

A ball is thrown upward and outward from the top edged of a 50 ft building it reaches its highest point 20 ft above and 10 ft out from the building. hmm vertex is as stated (10,50+20); vertex form of the equation is then y = a(x-10)^2 + 70 ; when x=0, then y=50 50 = a(-10)^2 + 70 -20 = a 100 -.2 = a therefore the equation would be: y = -.2(x-10)^2 + 70, so when does y=0? 0 = -.2(x-10)^2 + 70 -70 = -.2(x-10)^2 350 = (x-10)^2 sqrt(350) = x-10 10 + sqrt(350) = x = 28.708... yep

OpenStudy (anonymous):

Lol glad you've got it :)

OpenStudy (amistre64):

;) just keeping the old grey matter nice and pliable, good job